OpenStudy (anonymous):

What are the zeroes of f(x) = x2 + 5x + 6?

OpenStudy (welshfella):

This will factor. You need 2 numbers a and b where a * b = +6 and a +b = +5 to give factors (x + a)(x + b)

OpenStudy (welshfella):

what do you think those numbers are?

OpenStudy (anonymous):

2 and 3

OpenStudy (welshfella):

right so we have (x + 2)(x + 3) = 0 can you find the zeroes now?

OpenStudy (anonymous):

-2 and -3

OpenStudy (welshfella):

right

OpenStudy (anonymous):

so is that it?

OpenStudy (anonymous):

so the answer is -2 and-3 or 2 and 3?

OpenStudy (welshfella):

OH sorry1 my bad thinking its -2 and -3. {-2, -3}

OpenStudy (anonymous):

k thx!

OpenStudy (welshfella):

yw